# Jet launching from the Kerr black hole magnetosphere: An electrogeodesic approach

### v1

@@ Introduction
$m\frac{\textrm{d}\varphi}{\textrm{d}\lambda}$ $=\frac{g_{tt}L+g_{t\varphi}E}{g_{tt}g_{\varphi\varphi}-g^{2}_{t\varphi}}-q\frac{g_{tt}A_{\varphi}-g_{t\varphi}A_{t}}{g_{tt}g_{\varphi\varphi}-g^{2}_{t\varphi}}$ (3.60)
When the magnetosphere is turned off, i.e. when $q=0$, one recovers the standard equations for the time and azimuthal motions of the neutral particles on the Kerr geometry. They are well known to be separable in term of the Mino time $\textrm{d}\tau=\textrm{d}\lambda/\Sigma$. The key challenge when $q\neq 0$ is that the new terms do not spoil this separability. Concretely, upon factorizing by the function $1/\Sigma$, they should appear as functions of the radius $r$ or of the polar angle $\theta$ only. Notice that this is not guaranteed by the presence of the Carter constant. However, it turns out to be the case. To see this, one can use that
@@
to show that
@@
$q\frac{g_{tt}A_{\varphi}-g_{t\varphi}A_{t}}{-\Delta\sin^{2}{\theta}}$ $=-q\frac{{\mathcal{P}}\cos{\theta}}{\Sigma\sin^{2}{\theta}}$ (3.63)
Upon an overall factor of $1/\Sigma$, both contributions only involve a function of the polar angle. This ensures that the equations for $\dot{t}$ and $\dot{\varphi}$ are indeed separable. They are explicitly given by
@@
$\Sigma m\frac{\textrm{d}\varphi}{\textrm{d}\lambda}$ $=\frac{L}{\sin^{2}{\theta}}+\frac{a}{\Delta}\left(2MrE-aL\right)+\frac{q{\mathcal{P}}\cos{\theta}}{\sin^{2}{\theta}}$ (3.65)
@@
We are now in position to provide the exact analytic solution to the electrogeodesic equation. For clarity, let us gather the four equations. Introducing the Mino time
$\textrm{d}\tau=\frac{\textrm{d}\lambda}{\Sigma(r(\lambda),\theta(\lambda))}$ (3.66)
@@
$m\frac{\textrm{d}\varphi}{\textrm{d}\tau}$ $=\frac{L+q{\mathcal{P}}\cos{\theta}}{\sin^{2}{\theta}}+\frac{a}{\Delta}\left(2MrE-aL\right)$ (3.67d)
@@
$\Theta(\theta)$ $={\mathcal{K}}-m^{2}a^{2}\cos^{2}{\theta}-\left(\frac{L+q{\mathcal{P}}\cos{\theta}}{\sin{\theta}}-aE\sin{\theta}\right)^{2}$ (3.69)
Let us introduce the specific conserved energy $\varepsilon$, the specific conserved angular momentum $\ell$, the reduced Carter constant $k$ and the charge-to-mass ratio $\kappa$ by
@@
$\frac{\textrm{d}\varphi}{\textrm{d}\tau}$ $=\frac{\ell+\kappa{\mathcal{P}}\cos{\theta}}{\sin^{2}{\theta}}+\frac{a}{\Delta}\left(2Mr\varepsilon-a\ell\right)$ (3.71d)
@@
$\tilde{\Theta}(\theta)$ $=k-a^{2}\cos^{2}{\theta}-\left(\frac{\ell+\kappa{\mathcal{P}}\cos{\theta}}{\sin{\theta}}-a\varepsilon\sin{\theta}\right)^{2}$ (3.73)
@@
$\tau-\tau_{0}$ $=\int^{r}_{r_{0}}\frac{\epsilon_{r}\textrm{d}\bar{r}}{\sqrt{{\mathcal{R}}(\bar{r})}}=\int^{\theta}_{\theta_{0}}\frac{\epsilon_{\theta}\textrm{d}\bar{\theta}}{\sqrt{\tilde{\Theta}(\bar{\theta})}}$ (3.74)
@@
$\varphi-\varphi_{0}$ $=a\int^{r}_{r_{0}}\frac{2M\bar{r}\varepsilon-a\ell}{\Delta(\bar{r})}\frac{\epsilon_{r}\textrm{d}\bar{r}}{\sqrt{{\mathcal{R}}(\bar{r})}}+\int^{\theta}_{\theta_{0}}\left[\frac{\ell+\kappa{\mathcal{P}}\cos{\bar{\theta}}}{\sin^{2}{\bar{\theta}}}\right]\frac{\epsilon_{\theta}\textrm{d}\bar{\theta}}{\sqrt{\tilde{\Theta}(\bar{\theta})}}$ (3.76)
@@
In this section, we use the exact solution to the electrogeodesic equation derived in the previous section to study the acceleration of charged particles near the poles and the formation of jets. We present the exact formula for the acceleration. Then, we classify the stable and unstable polar equilibrium positions which define the opening angle of the jet at its basis. We further derive the expression for the azimuthal speed experienced by the charged particles in this jet. Finally, we compute the electromagnetic gravitational redshift measured by an asymptotic Carter observer and determine the condition under which the ejected charged particle is blueshifted. To our knowledge, the only other investigations of the motion of charged particles in the Kerr monopole have been discussed in [55] and more recently in [50].
We have seen that the electrogeodesics satisfies general conditions given by (3.16), (3.20) and (3.22). Applied to the magnetic monopole magnetosphere given by [cf. Eq. (2.25)]
$A_{\mu}\textrm{d}x^{\mu}=\frac{{\mathcal{P}}\cos{\theta}}{\Sigma}\left(a\textrm{d}t-(r^{2}+a^{2})\textrm{d}\varphi\right)$ (4.1)
@@
With this solution at end, one can obtain an analytic formula for the 4-acceleration $a^{\mu}$ of the charged particles, which reads
@@
It is easier to perform the computation in the Carter tetrad $(e_{\alpha})$, as given by Eq. (2.14). The components of the particle’s 4-velocity with respect to the coordinate frame are $u^{\alpha}=\Sigma^{-1}\textrm{d}x^{\alpha}/\textrm{d}\tau$ with $\textrm{d}x^{\alpha}/\textrm{d}\tau$ given by the electrogeodesic system (3.71). Transforming to the Carter tetrad via the system (2.14), we get
@@
Using the components (2.24) of the Faraday tensor in the Carter tetrad, Eq. (4.5) leads to the particle’s 4-acceleration in the Carter frame [^5] See Notebook 2 in App. A for the computation.:
@@
Consider first the equatorial plane, i.e. $\theta=\pi/2$. The components $a^{0}$ and $a^{1}$ of the 4-acceleration are proportional to $\cos{\theta}$, such that they vanish identically at the equator:
$a^{0}(r,\pi/2)=a^{1}(r,\pi/2)=0$ (4.9)
@@
$a^{3}(r,\pi/2)$ $=-\epsilon_{\theta}\kappa{\mathcal{P}}\frac{\sqrt{k-(\ell-a\varepsilon)^{2}}}{r^{1/2}}$ (4.11)
@@
On the contrary, near the poles, at $\theta=0$ (or equivalently $\theta=\pi$), the time and radial components of the acceleration $(a^{0},a^{1})$ take their maximal value, which suggests that charged particles are ejected preferentially near the poles. Let us also point that thanks to condition (4.3), the term $(r^{2}+a^{2})\varepsilon-a\ell$ in the radial acceleration (4.7b) cannot vanish.
@@
Recall that $k\geq 0$; hence we must have a not too large Carter constant: $0\leq k\leq a^{2}\cos^{2}\theta$. At this radius, the electrogeodesics reduces to a free-fall motion. Properly characterizing the role of this boundary in the magnetosphere will require more work which will be presented elsewhere. However, we already anticipate that this locus could play a role in allowing spherical non-accelerated trajectories.
Finally, it is interesting to obtain the behavior of the acceleration at large $r$ which reads
@@
while the the norm behaves as
@@
This concludes our discussion on the acceleration. It is crucial to notice that the present analysis has been made possible thanks to the exact solution of the electrogeodesic motion. As we can see, the expression for the acceleration experienced by the test charged particles already reveals several interesting features of the magnetosphere and the motion it induces on the particles. The key point is that even in this vacuum magnetosphere, the charged particles are naturally driven towards the poles and preferentially ejected near the poles.
@@
We can now study how to characterize the jet. To this end, it is useful to identify the equilibrium positions of the polar motion and the stable versus unstable nature. Since the idealized jet corresponds to the charged particles escaping at infinity along the rotation axis, we first focus on characterizing the conditions for the pole to be a stable position. As we shall see, the presence of the magnetosphere introduces again two regimes depending on the energy of the charged particles. Then, we further discuss the case of the equatorial plane.
Stable versus unstable polar equilibrium positons
@@
$\ell=-\kappa{\mathcal{P}}.$ (4.15)
to cancel the divergence in the potential. Notice that in the neutral case, one recovers the standard result that $\ell=0$. Moreover, for the value (4.15) of $\ell$ the general condition (4.3), i.e. $\varepsilon>\Omega_{\rm C}\ell$, reduces to condition (4.4), namely $\varepsilon>-\Omega_{\rm C}\kappa{\mathcal{P}}$.
@@
By definition, the equilibrium polar positions for the charged particles satisfy
$V(\theta)=0\qquad V^{\prime}(\theta)=0$ (4.20)
and it is stable by $V^{\prime\prime}(\theta)>0$ and unstable if $V^{\prime\prime}(\theta)<0$. The effective potential $V(\theta)$ is depicted in Figs. 12.
Let us first consider the north pole $\theta=0$ which corresponds to the idealized jet, i.e. to charged particles escaping to infinity along the rotation axis. In this case, one finds
$V(0)$ $=a^{2},\qquad V^{\prime}(0)=0,\qquad V^{\prime\prime}(0)=2a^{2}(\varepsilon-\beta/2-1)(\varepsilon-\beta/2+1)$ (4.21)
It follows that $\theta=0$ is a stable constant latitude position, i.e. that $\textrm{d}\theta/\textrm{d}\tau=0$ and $V^{\prime\prime}(0)>0$, if and only if, the electrogeodesic $(\varepsilon,\ell,k)$ satisfies
@@
For the second condition, we have taken into account that $a\Omega_{\rm C}=a^{2}/(r^{2}+a^{2})$ always fulfils $a\Omega_{\rm C}<1/2$, with the upper bound $1/2$ approached only near the event horizon of an extremal ($a=M$) Kerr black hole, so that $\varepsilon>a\Omega_{C}\beta$ is actually implied by $\varepsilon>{\beta}/{2}+1$.
It follows that there are two distinct families of electrogeodesics defined by the above thresholds which corresponds to the idealized jet. The expression of $V^{\prime}(\theta)$ is
$V^{\prime}(\theta)=\frac{2a^{2}}{\sin^{2}\theta}\left\{2\beta^{2}\tan\frac{\theta}{2}+\sin\theta\left[\left((\varepsilon^{2}-1)\cos\theta-\varepsilon\beta\right)\sin^{2}\theta-\beta^{2}\right]\right\}$ (4.24)
but note that the potential $V(\theta)$ is meaningful only for $\ell=-\kappa{\mathcal{P}}$.
@@
Figure 1: Effective potential $V(\theta)$ governing the latitudal motion for $\ell=-\kappa{\mathcal{P}}$. The plots are performed for $\beta=8$ and for various values of $\varepsilon$ ensuring a stable equilibrium at $\theta=0$, according to Eq. (4.22), i.e. $\varepsilon<3$ (left panel) or $\varepsilon>5$ (right panel). Notice that in all cases, $\theta=0$ is a (local) minimum of $V$ and that the marginal cases $\varepsilon=3$ and $\varepsilon=5$ are included in the plots.
Figure 2: Effective potential $V(\theta)$ governing the latitudal motion for $\ell=-\kappa{\mathcal{P}}$ and $\beta=8$. Contrary to Fig. 1, the chosen values of $\varepsilon$ are such that the equilibrium position at $\theta=0$ is unstable: it corresponds to a local maximum of $V(\theta)$. According to Eq. (4.22) with $\beta=8$, these values are in the range $3<\varepsilon<5$ (the marginal cases $\varepsilon=3$ and $\varepsilon=5$ are included in the plots). Notice that a stable equilibrium position occurs at $\theta=\theta_{*}\in(0,\pi/2)$.
To conclude this section, let us further comment on the parameter $\beta$. First, notice that for an electron [^7] In terms of SI units, one has actually $\kappa=\sqrt{\mu_{0}/(4\pi)}\,q/m$, so that, for an electron, $\kappa_{\rm e}=-5.56\;10^{7}\;{\rm m}^{1/2}{\rm kg}^{-1/2}$, which yields $\kappa_{\rm e}=-2.04\;10^{21}$ in geometrical units., $\kappa=\kappa_{\rm e}=-2.04\;10^{21}$, while for a proton, $\kappa=\kappa_{\rm p}=1.11\;10^{18}$. These huge values reflect the large dominance of the electromagnetic interaction over the gravitational one for elementary particles. On the contrary, $\mathcal{P}$} must obey $|{\mathcal{P}}|\ll M$, in order for the electromagnetic field to not significantly modify the spacetime metric, which is assumed to be the Kerr metric; otherwise one would have to consider the Kerr-Newman metric. For values of $a\sim M$, this is equivalent to demand $|{\mathcal{P}}|/a\ll 1$. It follows that $\beta$ can take values of order one or even much higher. For instance, for electrons, for which $|\kappa|\sim 10^{21}$, $\beta$ can be as large as $10^{10}$ for $|{\mathcal{P}}|=10^{-10}M$. This confirms that the model can account for large magnetization without breaking the test field assumption.
@@
$\frac{\textrm{d}\varphi}{\textrm{d}\tau}$ $=\frac{a}{\Delta}\left(2Mr\varepsilon+a\kappa{\mathcal{P}}\right)+\frac{\kappa{\mathcal{P}}\left(\cos{\theta}-1\right)}{\sin^{2}{\theta}}$ (4.26)
@@
$\Omega(r,\theta)=\frac{\textrm{d}\varphi}{\textrm{d}t}=\frac{\left(2Mr\varepsilon+a\kappa{\mathcal{P}}\right)a\sin^{2}{\theta}+\kappa{\mathcal{P}}\left(\cos{\theta}-1\right)\Delta}{\left[(r^{2}+a^{2})^{2}\varepsilon+2\kappa aM{\mathcal{P}}r+a\Delta\left(\kappa{\mathcal{P}}\cos{\theta}-a\varepsilon\sin^{2}{\theta}\right)\right]\sin^{2}{\theta}}$ (4.27)
As expected, this angular velocity is not defined for particles located at the poles, i.e. at $\theta=0$. However, this formula provides the angular velocity of the particle for $\theta\in\;]0,\pi/2]$.
Let us now evaluate its expression in the large $r$ limit. Before proceeding, notice that in the absence of a magnetosphere, i.e. ${\mathcal{P}}=0$, one recovers the standard gravitational frame-dragging effect of the Kerr black hole
$\Omega^{\text{K}}_{\infty}(r,\theta)\sim\frac{2aM}{r^{3}}$ (4.28)
such that for neutral particles sufficiently close to the black hole, the frame-dragging forces them to rotates with the hole.
The magnetic monopole dramatically changes this effect. When the magnetosphere is turned on, the large distance behavior is dominated by the electromagnetic interaction such that the azimuthal speed experienced by charged particle in our magnetized Kerr (MK) reads
$\Omega^{\text{MK}}_{\infty}(r,\theta)\sim-\frac{\kappa{\mathcal{P}}}{\varepsilon r^{2}}\frac{(1-\cos{\theta})}{\sin^{2}{\theta}}$ (4.29)
We observer that for ${\mathcal{P}}>0$, the angular velocity has the opposite sign of $\kappa$, i.e. the magnetosphere forces the positive charges to have a retrograde motion while negatve charge have a prograde motion. For a collimated jet, i.e. for a small opening angle $\theta_{m}$, it reduces to
$\Omega^{\text{MK}}_{\infty}(r,\theta)\sim-\frac{\kappa{\mathcal{P}}}{2\varepsilon r^{2}}\left(1+\frac{\theta_{m}^{2}}{4}\right)+{\mathcal{O}}(\theta_{m}^{4})$ (4.30)
Remarkably, it depends solely on the magnetic monopole charge $\mathcal{P}$}, the mass to charge ratio $\kappa$ and the energy of the particle $\varepsilon$. At the equator $\theta=\pi/2$, one finds
$\Omega^{\text{MK}}_{\infty}(r,\theta)\sim-\frac{\kappa{\mathcal{P}}}{\varepsilon r^{2}}$ (4.31)
To conclude, the magnetosphere induces a new magnetic frame-dragging effect which completely dominates its usual gravitational frame dragging counterpart. Moreover, one notice that while the gravitational frame dragging effect decays as $1/r^{3}$ at large distance, the magnetic frame dragging decays only as $1/r^{2}$, showing that it remains effective on much larger distance away from the black hole.

### v2

@@ Introduction
$m\frac{\textrm{d}\varphi}{\textrm{d}\lambda}$ $=\frac{g_{tt}L+g_{t\varphi}E}{g_{tt}g_{\varphi\varphi}-g^{2}_{t\varphi}}-q\frac{g_{tt}A_{\varphi}-g_{t\varphi}A_{t}}{g_{tt}g_{\varphi\varphi}-g^{2}_{t\varphi}}.$ (3.59)
When the magnetosphere is turned off, i.e. when $q=0$, one recovers the standard equations for the time and azimuthal motions of the neutral particles on the Kerr geometry. When $q\neq 0$, the equations inherit new terms. Using that
@@
one can show that these new terms read
@@
$q\frac{g_{tt}A_{\varphi}-g_{t\varphi}A_{t}}{-\Delta\sin^{2}{\theta}}$ $=-q\frac{{\mathcal{P}}\cos{\theta}}{\Sigma\sin^{2}{\theta}}.$ (3.62)
Upon an overall factor of $1/\Sigma$, both contributions only involve a function of the polar angle. Upon using the radial and polar equations, this ensures that the equations for $\dot{t}$ and $\dot{\varphi}$ are indeed separable. They are explicitly given by
@@
$\Sigma m\frac{\textrm{d}\varphi}{\textrm{d}\lambda}$ $=\frac{L}{\sin^{2}{\theta}}+\frac{a}{\Delta}\left(2MrE-aL\right)+\frac{q{\mathcal{P}}\cos{\theta}}{\sin^{2}{\theta}}.$ (3.64)
@@
We are now in position to present the exact analytic solution to the electrogeodesic equation in the Kerr magnetic monopole. For clarity, let us gather the four equations. Introducing the Mino time
$\textrm{d}\tau=\frac{\textrm{d}\lambda}{\Sigma(r(\lambda),\theta(\lambda))},$ (3.65)
@@
$m\frac{\textrm{d}\varphi}{\textrm{d}\tau}$ $=\frac{L+q{\mathcal{P}}\cos{\theta}}{\sin^{2}{\theta}}+\frac{a}{\Delta}\left(2MrE-aL\right),$ (3.66d)
@@
$\Theta(\theta)$ $={\mathcal{K}}-m^{2}a^{2}\cos^{2}{\theta}-\left(\frac{L+q{\mathcal{P}}\cos{\theta}}{\sin{\theta}}-aE\sin{\theta}\right)^{2}.$ (3.68)
As expected, this system of equations corresponds to the electrically neutral subsector of the equations derived and studied by Hackmann and Xu in [56].
For later purposes, it will be useful to introduce the specific conserved energy $\varepsilon$, the specific conserved angular momentum $\ell$, the reduced Carter constant $k$ and the charge-to-mass ratio $\kappa$ by
@@
$\frac{\textrm{d}\varphi}{\textrm{d}\tau}$ $=\frac{\ell+\kappa{\mathcal{P}}\cos{\theta}}{\sin^{2}{\theta}}+\frac{a}{\Delta}\left(2Mr\varepsilon-a\ell\right),$ (3.70d)
@@
$\tilde{\Theta}(\theta)$ $=k-a^{2}\cos^{2}{\theta}-\left(\frac{\ell+\kappa{\mathcal{P}}\cos{\theta}}{\sin{\theta}}-a\varepsilon\sin{\theta}\right)^{2}.$ (3.72)
@@
$\tau-\tau_{0}$ $=\int^{r}_{r_{0}}\frac{\epsilon_{r}\textrm{d}\bar{r}}{\sqrt{{\mathcal{R}}(\bar{r})}}=\int^{\theta}_{\theta_{0}}\frac{\epsilon_{\theta}\textrm{d}\bar{\theta}}{\sqrt{\tilde{\Theta}(\bar{\theta})}}.$ (3.73)
@@
$\varphi-\varphi_{0}$ $=a\int^{r}_{r_{0}}\frac{2M\bar{r}\varepsilon-a\ell}{\Delta(\bar{r})}\frac{\epsilon_{r}\textrm{d}\bar{r}}{\sqrt{{\mathcal{R}}(\bar{r})}}+\int^{\theta}_{\theta_{0}}\left[\frac{\ell+\kappa{\mathcal{P}}\cos{\bar{\theta}}}{\sin^{2}{\bar{\theta}}}\right]\frac{\epsilon_{\theta}\textrm{d}\bar{\theta}}{\sqrt{\tilde{\Theta}(\bar{\theta})}}.$ (3.75)
@@
In this section, we use the exact solution to the electrogeodesic equation derived in the previous section to study the acceleration of charged particles near the poles and the formation of jets. We present the exact formula for the 4-acceleration. Then, we classify the stable and unstable polar equilibrium positions. We further derive the expression for the azimuthal speed experienced by the charged particles in this jet, which reveals the role of the magnetic frame-dragging effect. Finally, we compute the electromagnetic gravitational redshift measured by an asymptotic Carter observer and determine the condition under which the ejected charged particle is blueshifted. To our knowledge, the only other analytical investigations of the motion of charged particles in the Kerr monopole have been discussed in [56] and more recently in [50].
We have seen that the electrogeodesics satisfies general conditions given by (3.16), (3.20) and (3.22). Applied to the magnetic monopole magnetosphere given by the 4-potential [cf. Eq. (2.24)]
$A=\frac{{\mathcal{P}}\cos{\theta}}{\Sigma}\left(a\textrm{d}t-(r^{2}+a^{2})\textrm{d}\varphi\right).$ (4.1)
@@
With this solution at end, one can obtain an analytic formula for the 4-acceleration $a^{\mu}$ of the charged particles. The latter is given by the Lorentz force formula:
@@
It is easier to perform the computation in the Carter tetrad $(e_{\alpha})$, as given by Eq. (2.13). The components of the particle’s 4-velocity with respect to the coordinate frame are $u^{\alpha}=\Sigma^{-1}\textrm{d}x^{\alpha}/\textrm{d}\tau$ with $\textrm{d}x^{\alpha}/\textrm{d}\tau$ given by the electrogeodesic system (3.70). Transforming to the Carter tetrad via the system (2.13), we get the following expression of the 4-velocity:
@@
Using the components (2.23) of the Faraday tensor in the Carter tetrad, Eq. (4.5) leads to the following components of the particle’s 4-acceleration in the Carter frame [^5] See Notebook 2 in App. A for the computation.:
@@
Consider first the equatorial plane, i.e. $\theta=\pi/2$. The components $a^{0}$ and $a^{1}$ of the 4-acceleration are proportional to $\cos{\theta}$, so that they identically vanish at the equator:
$a^{0}(r,\pi/2)=a^{1}(r,\pi/2)=0.$ (4.9)
@@
$a^{3}(r,\pi/2)$ $=-\epsilon_{\theta}\kappa{\mathcal{P}}\frac{\sqrt{k-(\ell-a\varepsilon)^{2}}}{r^{1/2}}.$ (4.11)
@@
On the contrary, near the poles, at $\theta=0$ (or equivalently $\theta=\pi$), the time and radial components of the acceleration $(a^{0},a^{1})$ take their maximal value, which suggests that charged particles are ejected preferentially near the poles. Notice that the time and radial components of the acceleration is directly proportional to (i) the spin $a$ of the black hole and (ii) the parameter $\kappa{\mathcal{P}}$ which encodes the presence of the magnetosphere, demonstrating the combined roles of rotation and the magnetic monopole to radially accelerate charged particles. In this regard, this property of the radial acceleration of charged particles in the magnetic monopole Kerr magnetosphere parallels the behavior of the Poynting flux in the Blandford-Znajek mechanism, which is also triggered by both the rotation of the black hole and the magnetic monopole charge of the magnetosphere. Finally, let us also point that thanks to condition (4.3), the term $(r^{2}+a^{2})\varepsilon-a\ell$ in the radial acceleration (4.7b) cannot vanish.
@@
Recall that $k\geq 0$; hence we must have a not too large Carter constant: $0\leq k\leq a^{2}\cos^{2}\theta$. At $r=r_{\ast}(\theta)$, the electrogeodesics reduces to a free-fall motion. Properly characterizing the role of this boundary in the magnetosphere will require more work, which will be presented elsewhere. However, we already anticipate that this locus could play a role in allowing spherical non-accelerated trajectories.
Finally, it is interesting to obtain the behavior of the acceleration at large $r$; it reads
@@
while the norm behaves as
@@
This concludes our discussion on the acceleration. It is crucial to notice that the present analysis has been made possible thanks to the exact solution of the electrogeodesic motion. As we can see, the expression for the acceleration experienced by the test charged particles already reveals several interesting features of the magnetosphere and the motion it induces on the particles. The key point is that even in this vacuum magnetosphere, the charged particles are naturally driven towards the poles and preferentially ejected near the poles, the ejection (i.e. the radial acceleration) being triggered both by the rotation and the magnetic monopole carried by the black hole.
@@
We now study how to characterize the jet. To this end, it is useful to identify the equilibrium positions of the polar motion and the stability of these positions. Since the aligned jet corresponds to charged particles escaping at infinity along the rotation axis, we first focus on characterizing the conditions for the rotation axis to be a stable position. As we shall see, the presence of the magnetosphere introduces two regimes depending on the energy of the charged particles. Then, we further discuss the case of the equatorial plane.
Stable versus unstable polar equilibrium positions
@@
$\ell=-\kappa{\mathcal{P}}$ (4.15)
in order to cancel the divergence in the potential. Notice that in the neutral case, one recovers the standard result that $\ell=0$ is required to reach the rotation axis. Moreover, for the value (4.15) of $\ell$, the general condition (4.3), i.e. $\varepsilon>\Omega_{\rm C}\ell$, reduces to condition (4.4), namely $\varepsilon>-\Omega_{\rm C}\kappa{\mathcal{P}}$.
@@
By definition, an equilibrium latitudinal position for the charged particles satisfies
$V(\theta)=k,\qquad V^{\prime}(\theta)=0$ (4.20)
and it is stable if $V^{\prime\prime}(\theta)>0$ and unstable if $V^{\prime\prime}(\theta)<0$. The effective potential $V(\theta)$ is depicted in Figs. 12.
Let us first consider the North pole $\theta=0$, which corresponds to the aligned jet, i.e. to charged particles escaping to infinity along the rotation axis. In this case, one finds
$V(0)$ $=a^{2},\qquad V^{\prime}(0)=0,\qquad V^{\prime\prime}(0)=2a^{2}(\varepsilon-\beta/2-1)(\varepsilon-\beta/2+1).$ (4.21)
It follows that $\theta=0$ is a stable constant latitudinal position, i.e. that $\textrm{d}\theta/\textrm{d}\tau=0$ and $V^{\prime\prime}(0)>0$, if and only if, the electrogeodesic $(\varepsilon,\ell,k)$ satisfies
@@
For the second condition, we have taken into account that $a\Omega_{\rm C}=a^{2}/(r^{2}+a^{2})$ always fulfills $a\Omega_{\rm C}<1/2$, with the upper bound $1/2$ approached only near the event horizon of an extremal ($a=M$) Kerr black hole, so that $\varepsilon>a\Omega_{C}\beta$ is actually implied by $\varepsilon>{\beta}/{2}+1$. It follows that there are two distinct families of electrogeodesics defined by the above thresholds, which correspond to such and aligned jet.
Beside $\theta=0$, other possible latitudinal equilibrium positions can be found from the zeros of $V^{\prime}(\theta)$, which reads, for $\ell=-\kappa{\mathcal{P}}$,
$V^{\prime}(\theta)=\frac{2a^{2}}{\sin^{2}\theta}\left\{2\beta^{2}\tan\frac{\theta}{2}+\sin\theta\left[\left((\varepsilon^{2}-1)\cos\theta-\varepsilon\beta\right)\sin^{2}\theta-\beta^{2}\right]\right\}.$ (4.24)
Figures 1 (right panel) and 2 reveal that stable latitudinal equilibria at $\theta=\theta_{*}\neq 0$ exist for some values of the parameters. They are also of interest as they may correspond to misaligned jets of charged particles. Their characterization is however more involved and shall be discussed in a future work.
@@
Figure 1: Effective potential $V(\theta)$ governing the latitudinal motion for $\ell=-\kappa{\mathcal{P}}$. The plots are performed for $\beta=8$ and for various values of $\varepsilon$ ensuring a stable equilibrium at $\theta=0$, according to Eq. (4.22), i.e. $\varepsilon<3$ (left panel) or $\varepsilon>5$ (right panel). Notice that in all cases, $\theta=0$ is a (local) minimum of $V$ and that the marginal cases $\varepsilon=3$ and $\varepsilon=5$ are included in the plots.
Figure 2: Effective potential $V(\theta)$ governing the latitudinal motion for $\ell=-\kappa{\mathcal{P}}$ and $\beta=8$. Contrary to Fig. 1, the chosen values of $\varepsilon$ are such that the equilibrium position at $\theta=0$ is unstable: it corresponds to a local maximum of $V(\theta)$. According to Eq. (4.22) with $\beta=8$, these values are in the range $3<\varepsilon<5$ (the marginal cases $\varepsilon=3$ and $\varepsilon=5$ are included in the plots). Notice that a stable equilibrium position occurs at $\theta=\theta_{*}\in(0,\pi/2)$.
To conclude this section, let us comment on the parameter $\beta$. First, notice that for an electron [^7] In terms of SI units, one has actually $\kappa=\sqrt{\mu_{0}/(4\pi)}\,q/m$, so that, for an electron, $\kappa_{\rm e}=-5.56\;10^{7}\;{\rm m}^{1/2}{\rm kg}^{-1/2}$, which yields $\kappa_{\rm e}=-2.04\;10^{21}$ in geometrical units., $\kappa=\kappa_{\rm e}=-2.04\;10^{21}$, while for a proton, $\kappa=\kappa_{\rm p}=1.11\;10^{18}$. These huge values reflect the large dominance of the electromagnetic interaction over the gravitational one for elementary particles. On the contrary, $\mathcal{P}$} must obey $|{\mathcal{P}}|\ll M$, in order for the electromagnetic field to not significantly alter the spacetime metric, which is assumed to be the Kerr metric; otherwise one would have to consider the Kerr-Newman metric. For values of $a\sim M$, this is equivalent to demand $|{\mathcal{P}}|/a\ll 1$. It follows that $\beta$ can take values of order one or even much higher. For instance, for electrons, for which $|\kappa|\sim 10^{21}$, $\beta$ can be as large as $10^{10}$ for $|{\mathcal{P}}|=10^{-10}M$. This confirms that the model can account for large magnetization without breaking the test field assumption.
@@
$\frac{\textrm{d}\varphi}{\textrm{d}\tau}$ $=\frac{a}{\Delta}\left(2Mr\varepsilon+a\kappa{\mathcal{P}}\right)+\frac{\kappa{\mathcal{P}}\left(\cos{\theta}-1\right)}{\sin^{2}{\theta}}.$ (4.26)
@@
$\Omega(r,\theta)=\frac{\textrm{d}\varphi}{\textrm{d}t}=\frac{\left(2Mr\varepsilon+a\kappa{\mathcal{P}}\right)a\sin^{2}{\theta}+\kappa{\mathcal{P}}\left(\cos{\theta}-1\right)\Delta}{\left[(r^{2}+a^{2})^{2}\varepsilon+2\kappa aM{\mathcal{P}}r+a\Delta\left(\kappa{\mathcal{P}}\cos{\theta}-a\varepsilon\sin^{2}{\theta}\right)\right]\sin^{2}{\theta}}.$ (4.27)
As expected, this angular velocity is not defined for particles located at the poles, i.e. at $\theta=0$. However, this formula provides the angular velocity of the particle for $\theta\in\;]0,\pi/2]$. In particular, if the jet is misaligned with the black hole rotation axis, i.e. $\theta\neq 0$, this angular velocity encodes the precession of the jet of charged particles. Such precession of a misaligned jet has been recently measured for the M87* supermassive black hole and reported in [61].
Let us now evaluate the expression of $\Omega(r,\theta)$ in the large $r$ limit. Before proceeding, notice that in the absence of a magnetosphere, i.e. ${\mathcal{P}}=0$, one recovers the standard gravitational frame-dragging effect of the Kerr black hole:
$\Omega^{\text{K}}_{\infty}(r,\theta)\sim\frac{2aM}{r^{3}}.$ (4.28)
The magnetic monopole dramatically changes this effect. When the magnetosphere is turned on, the large distance behavior is dominated by the electromagnetic interaction: the azimuthal speed experienced by charged particle in our magnetized Kerr (MK) becomes
$\Omega^{\text{MK}}_{\infty}(r,\theta)\sim-\frac{\kappa{\mathcal{P}}}{\varepsilon r^{2}}\frac{(1-\cos{\theta})}{\sin^{2}{\theta}}.$ (4.29)
We observe that for ${\mathcal{P}}>0$, the angular velocity has a sign opposite to that of $\kappa$, i.e. the magnetosphere forces positive charges to have a retrograde motion, while negative charges have a prograde motion. For a collimated jet, i.e. for a small opening angle $\theta_{\rm m}$, it reduces to
$\Omega^{\text{MK}}_{\infty}(r,\theta_{\rm m})\sim-\frac{\kappa{\mathcal{P}}}{2\varepsilon r^{2}}\left(1+\frac{\theta_{\rm m}^{2}}{4}\right)+{\mathcal{O}}(\theta_{\rm m}^{4}).$ (4.30)
Remarkably, it depends solely on the magnetic monopole charge $\mathcal{P}$}, the charge-to-mass ratio $\kappa$ and the specific energy of the particle $\varepsilon$. At the equator $\theta=\pi/2$, Eq. (4.29) yields
$\Omega^{\text{MK}}_{\infty}(r,\theta)\sim-\frac{\kappa{\mathcal{P}}}{\varepsilon r^{2}}.$ (4.31)
To conclude, the magnetosphere induces a new magnetic frame-dragging effect which completely dominates the standard gravitational frame dragging for electrogeodesics with $\ell=-\kappa{\mathcal{P}}$. Moreover, one notices that while the gravitational frame dragging effect decays as $1/r^{3}$ at large distance, the magnetic frame dragging decays only as $1/r^{2}$, showing that it remains effective on much larger distances away from the black hole.